# Image is All You Need: Towards Efficient and Effective Large Language Model-Based Recommender Systems

### v1

@@ 3.1. Mapping Item Images to LLM
Figure 3. Overall framework of I-LLMRec. User-interacted item images are mapped into the LLM through an adaptor, which bridges the image and language spaces. To ensure alignment between two spaces, the adaptor is optimized via the ILA module. Furthermore, the recommendation process is formulated as a retrieval task via the IRE module.
3.1.2. Designing a prompt to represent a user’s item interaction history.
To represent a user’s item interaction history with visual features, we carefully design a prompt, enabling the LLM to comprehensively understand the item semantics. Specifically, for each item $i$ in a user’s item interaction history, we represent it in the prompt as follows:
@@
3.2. Image-LLM Alignment (ILA) Module
To make the LLM effectively capture user preferences from item images, it is crucial to align visual features with the language space in a meaningful way, requiring adequate supervision of the adaptor network $M$. To achieve this, we optimize the adaptor network $M$ by predicting the next item properties based on the user interaction within a structured prompt. This approach ensures that the LLM learns to associate visual features with relevant textual information, thereby improving its ability to understand user preferences in the context of recommendation.
Specifically, we craft a prompt in an "input"-"target output" format. Here, the input consists of the user interaction prompt, followed by a question regarding the next item (e.g., What brand is this user likely to consume for the next item?). The target output corresponds to the answer to that question (e.g., The brand likely to be consumed is ‘WOLT’). To guide the LLM in considering various properties of the next item, we incorporate multiple properties of the next item, including brand, category, title, and description. For each property, we design five different question templates (refer to Appendix I for complete templates). During each training step, we randomly select one of 20 possible question templates (4 properties × 5 templates) and train $M$ to generate the corresponding output. The training objective is formulated as follows:
[…]
where $\theta$ is the parameter of the LLM that is frozen, $y$ is the target output, $y_{k}$ is the $k$-th token of $y$, and $x$ is the input.
3.2.1. Discussion regarding LLM fine-tuning
A further benefit of I-LLMRec is that LLM fine-tuning is not mandatory, unlike the existing LLM-based recommender systems (Lin et al., 2024; Bao et al., 2023; Tan et al., 2024; Zheng et al., 2024; Liu et al., 2024a) that generally rely on costly fine-tuning of the LLM. Instead, we harness the LLM’s intrinsic parametric knowledge to understand the user preferences. Even though I-LLMRec could be potentially improved by fine-tuning the LLM (See Appendix E), we opt to keep the LLM frozen to preserve the LLM’s intrinsic knowledge and improve the training efficiency.
3.3. Image-based REtrieval (IRE) Module
We found that existing studies on LLM-based recommender systems often face challenges in providing reliable and efficient recommendations. This is because they typically recommend items by predicting the next item title (Kim et al., 2024a; Lin et al., 2024; Tan et al., 2024) or item tokens allocated as out-of-vocabulary (Wei et al., 2024; Li et al., 2023b; Zhu et al., 2024). Regarding reliable recommendation, the title prediction approach cannot guarantee that the recommended items exist within the item corpus (i.e., lack of item grounding), which sometimes may lead to recommending non-existent items. Regarding efficient recommendation, title prediction relies on computationally intensive beam search to recommend multiple items, while the item token prediction approach requires an oversized LLM, as expanding the item pool necessitates extending the LLM vocabulary set, which hinders scalability. To overcome these challenges, we propose the Image-based REtrieval (IRE) module, which formulates the recommendation task as a retrieval task by leveraging the images to directly retrieve relevant items from the item pool. In the following, we describe how to obtain an LLM-guided user representation for retrieval, and define a training objective that retrieves relevant items using the visual features.
3.3.1. LLM-guided user representation
[…]
3.3.2. Training objective for retrieving relevant items
[…]
(4) $o_{u}^{\textsf{Img}}=F^{\textsf{Img}}_{u}(h(\left[\texttt{REC}\right])),o_{i}^{\textsf{Img}}=F^{\textsf{Img}}_{i}(v_{i_{|\mathcal{S}_{u}|}})$
where $F^{\textsf{Img}}_{u}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{s}}$, $F^{\textsf{Img}}_{i}:\mathbb{R}^{d_{v}}\rightarrow\mathbb{R}^{d_{s}}$ are the projectors, and $d_{s}$ is the feature dimension of shared recommendation space. $o_{u}^{\textsf{Img}}$ and $o_{i}^{\textsf{Img}}$ denote the projected representations of user $u$ and item $i$ in terms of visual features, respectively. To optimize the retrieval task, we use the binary cross-entropy loss as the training objective:
(5) $\begin{split}\mathcal{L}_{\text{IRE}}^{\textsf{Img}}&=-\sum_{u\in\mathcal{U}}\left[log(\sigma(r_{u,i}^{\textsf{Img}}))+log(1-\sigma(r_{u,i^{-}}^{\textsf{Img}}))\right]\end{split}$
[…]
3.3.3. Extension to multiple feature types.
Our retrieval-based recommendation approach can seamlessly incorporate multiple item feature types (e.g., ID-based embeddings and textual features) alongside visual features. Specifically, for a given item feature type denoted as *, we can simply integrate it by introducing two additional projectors, $F^{*}_{u}$ and $F^{*}_{i}$:
(6) $o_{u}^{*}=F^{*}_{u}(h(\left[\texttt{REC}\right])),o_{i}^{*}=F^{*}_{i}(\mathbf{IF}^{*})$
where $\mathbf{IF}^{*}$ represents the item feature corresponding to feature type * (e.g., $v_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{Img}}$ if * is the image type Img), and $o_{u}^{*}$ and $o_{i}^{*}$ are the projected representations of user $u$ and item $i$ in terms of * feature type, respectively. Using $o_{u}^{*}$ and $o_{i}^{*}$, we can then compute the binary cross-entropy loss $\mathcal{L}_{\text{IRE}}^{*}$ following Equation 5.
Among various item feature types, we extract the ID-based item embeddings from a pretrained collaborative filtering (CF) model (i.e., SASRec (Kang and McAuley, 2018)), thanks to its effectiveness in general recommendation tasks, especially for warm items (Kim et al., 2024a) [^7] In Figure 7, we show that the utilization of ID-based item embeddings leads to more recommendations of warm items.. Specifically, we incorporate ID-based item embeddings (i.e., $c_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{CF}}$) in addition to the visual features (i.e., $v_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{Img}}$). Furthermore, as item descriptions are readily available in real-world scenarios (Zheng et al., 2024), we can easily incorporate textual features $t$ extracted from full descriptions to further improve overall performance (i.e., $t_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{Text}}$). The analysis of extension to various features is provided in Section 4.4.
3.4. Model Training
[…]
(7) $\begin{split}\mathcal{L}_{final}=\mathcal{L}_{\text{ILA}}+\mathcal{L}_{\text{IRE}}^{\textsf{Img}}+\mathcal{L}_{\text{IRE}}^{\textsf{CF}}+\mathcal{L}_{\text{IRE}}^{\textsf{Text}}\end{split}$
3.5. Inference
For retrieval-based inference, we compute the affinity scores $r_{u,i}^{\textsf{Img}}$, $r_{u,i}^{\textsf{CF}}$, and $r_{u,i}^{\textsf{Text}}$ using the scoring function $\mathcal{T}$, in terms of visual, CF, and textual features, respectively. Specifically, the top-$k$ relevant items for $i_{|\mathcal{S}_{u}|+1}$ is computed as follows:
(8) $rec_{u}^{k}=\textsf{Top-k}(r_{u,i}^{\textsf{Img}}+r_{u,i}^{\textsf{CF}}+r_{u,i}^{\textsf{Text}}),\quad\forall i\in\mathcal{I}$
where $rec_{u}^{k}$ is the set of recommended $k$ items for user $u$, and Top-k is a function that extracts items with the highest top-$k$ scores. Note that while the affinity scores across different features can be aggregated through various approaches, such as weight summation or product, we opt for a simpler summation.
Table 1. Performance Comparison. The best performance is denoted in bold text, while the second-best among baselines is underlined. $\mathbf{A}$: Attributed-based Representation, $\mathbf{CF}$: CF-based Representation (i.e., the CF item embedding is projected into the LLM), $\mathbf{D}$: Description-based Representation, $\mathbf{I}$: Image-based Representation.
Dataset Metric Collaborative Filtering LLM Image-aware LLM
GRU4Rec VBPR BERT4Rec SASRec TALLRec ($\mathbf{A}$) A-LLMRec ($\mathbf{CF}$) TRSR ($\mathbf{D}$) UniMP ($\mathbf{I}$) I-LLMRec ($\mathbf{I}$)+$\mathbf{D}$ I-LLMRec ($\mathbf{I}$)

### v2

@@ 3.1 Mapping Item Images to an LLM
Figure 2: Overall framework of I-LLMRec. User-interacted item images are mapped into the LLM through an adaptor, which bridges the image and language spaces. To ensure alignment between two spaces, the adaptor is optimized via the RISA module. Furthermore, the recommendation process is formulated as a retrieval task via the RERI module.
Designing a prompt to represent a user’s item interaction history. To represent a user’s item interaction history with visual features, we carefully design a prompt, enabling the LLM to comprehensively understand the item semantics. Specifically, for each item $i$ in a user’s item interaction history, we represent it in the prompt as $\mathbf{P}_{i}=\left\{\textrm{Title}:\textsf{ITEM\_TITLE},\textrm{Visual}\;\textrm{Representation}:\left[\textsf{VISUAL}\right]\right\}$, where $\mathbf{P}_{i}$ is the representation of item $i$, ITEM_TITLE is the title of the item $i$, and $\left[\textsf{VISUAL}\right]$ is the placeholder of item $i$’s visual feature $\bar{v}_{i}$. More precisely, before being forwarded to the LLM’s transformer layer, $\left[\textsf{VISUAL}\right]$ is replaced with $\bar{v}_{i}$, while the natural language in $\mathbf{P}_{i}$ is converted into word embeddings. Note that following prior studies (Bao et al., 2023; Li et al., 2023a), we use the item title instead of the item’s numerical ID. By enumerating $\mathbf{P}_{i}$ over the item interaction history of user $u$ (i.e., $\mathbf{P}_{i_{1}},\mathbf{P}_{i_{2}},...,\mathbf{P}_{i_{|\mathcal{S}_{u}|-1}}]$), we represent the user interaction based on item images where each image is expressed using a single token, allowing the LLM to efficiently and effectively capture the user preference.
3.2 Recommendation-Oriented Image–LLM Semantic Alignment (RISA) Module
To make the LLM effectively capture user preferences from item images, it is crucial to align visual features with the language space in a meaningful way, requiring adequate supervision of the adaptor network $M$. To achieve this, we optimize the adaptor network $M$ by predicting the next item properties based on the user interaction within a structured prompt, thereby guiding the LLMs to capture user preferences in the context of recommendation through the visual features.
Specifically, we craft a prompt in an "input"-"target output" format. Here, the input consists of the user interaction prompt, followed by a question regarding the next item (e.g., What brand is this user likely to consume for the next item?). The target output corresponds to the answer to that question (e.g., The brand likely to be consumed is ‘WOLT’). To guide the LLM in considering various properties of the next item, we incorporate multiple properties of the next item, including brand, category, title, and description. For each property, we design five different question templates (refer to Appendix D.1 for complete templates). During each training step, we randomly select one of 20 possible question templates (4 properties × 5 templates) and train $M$ to generate the corresponding output. The training objective is formulated as $\mathcal{L}_{\text{RISA}}=\underset{M}{\textrm{max}}\sum_{k=1}^{|y|}log(P_{\theta,M}(y_{k}|x,y_{<k}))$, where $\theta$ is the parameter of the LLM that is frozen, $y$ is the target output, $y_{k}$ is the $k$-th token of $y$, and $x$ is the input. Note that $x$ incorporates the image features of items consumed by users, which are supervised under $\mathcal{L}_{\text{RISA}}$. Please refer to Appendix D.2 for further discussion of LLM fine-tuning.
3.3 REtrieval-based Recommendation via Image features (RERI) module
We found that existing studies on LLM-based recommender systems often face challenges in providing reliable and efficient recommendations since they typically recommend items by predicting the next item title (Kim et al., 2024b; Lin et al., 2024b; Tan et al., 2024) or item tokens allocated as out-of-vocabulary (Wei et al., 2024; Li et al., 2023a; Zhu et al., 2024). Regarding reliable recommendation, the title prediction approach cannot guarantee that the recommended items exist within the item corpus, being likely to recommend non-existent items. Regarding efficient recommendation, title prediction relies on computationally intensive beam search to recommend multiple items, while the item token prediction approach requires an oversized LLM, as expanding the item pool necessitates extending the LLM vocabulary set, which hinders scalability. To overcome these challenges, we propose the REtrieval-based Recommendation via Image features (RERI) module, which formulates the recommendation task as a retrieval task by leveraging the images to directly retrieve relevant items from the item pool. In the following, we describe how to obtain an LLM-guided user representation for retrieval, and define a training objective that retrieves relevant items.
LLM-guided user representation. To obtain the LLM-guided user representation containing user preference, for each user $u$, we append an instruction prompt, i.e., Generate a recommendation token for the next item to be consumed, after the user interaction prompt $\mathbf{P}_{i_{1}},\mathbf{P}_{i_{2}},...,\mathbf{P}_{i_{|\mathcal{S}_{u}|-1}}]$. This instruction prompt helps generate a recommendation token, guiding the model towards item retrieval-based recommendations instead of item title generation-based recommendation. At the end of this prompt, we append a learnable token $\left[\texttt{REC}\right]$ that aggregates a user’s item interaction history and instruction-related information, leveraging the LLM’s contextual reasoning capabilities. Specifically, we utilize the last hidden state associated with the $\left[\texttt{REC}\right]$ token (i.e., hh($\left[\texttt{REC}\right]$)) to obtain the aggregated information of user preference for retrieval.
Training objective for retrieving relevant items. Given that hh($\left[\texttt{REC}\right]$) represents the user’s preference contained in the interaction history of user $u$, i.e., $i_{1},i_{2},...,i_{|\mathcal{S}_{u}|-1}$], we formulate the task of retrieving the target item $i_{|\mathcal{S}_{u}|}$ based on a scoring function $\mathcal{T}$, unlike previous studies relying on a generative task (i.e., item title generation). To this end, we use the visual feature of items as the item representations and compute the affinity score between those and hh($\left[\texttt{REC}\right]$). However, as hh($\left[\texttt{REC}\right]$) and the visual features lie in the different spaces, we employ projectors for each, ensuring that they are mapped into a shared recommendation space. We denote the projectors and their outputs as $o_{u}^{\textsf{Img}}=F^{\textsf{Img}}_{u}(h(\left[\texttt{REC}\right])),o_{i}^{\textsf{Img}}=F^{\textsf{Img}}_{i}(v_{i_{|\mathcal{S}_{u}|}})$ , where $F^{\textsf{Img}}_{u}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{s}}$, $F^{\textsf{Img}}_{i}:\mathbb{R}^{d_{v}}\rightarrow\mathbb{R}^{d_{s}}$ are the projectors, and $d_{s}$ is the feature dimension of shared recommendation space. $o_{u}^{\textsf{Img}}$ and $o_{i}^{\textsf{Img}}$ denote the projected representations of user $u$ and item $i$ in terms of visual features, respectively. To optimize the retrieval task, we use the binary cross-entropy loss as follows:
$\begin{split}\mathcal{L}_{\text{RERI}}^{\textsf{Img}}&=-\sum_{u\in\mathcal{U}}\left[log(\sigma(r_{u,i}^{\textsf{Img}}))+log(1-\sigma(r_{u,i^{-}}^{\textsf{Img}}))\right]\end{split}$ (1)
@@ 3.3 REtrieval-based Recommendation via Image features (RERI) module
Extension to multiple feature types. Our retrieval-based recommendation approach can seamlessly incorporate multiple item feature types (e.g., ID-based embeddings and textual features) alongside visual features. Specifically, for a given item feature type denoted as *, we can simply integrate it by introducing two additional projectors, $F^{*}_{u}$ and $F^{*}_{i}$, such that $o_{u}^{*}=F^{*}_{u}(h(\left[\texttt{REC}\right])),o_{i}^{*}=F^{*}_{i}(\mathbf{IF}^{*})$, where $\mathbf{IF}^{*}$ represents the item feature corresponding to feature type * (e.g., $v_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{Img}}$ if * is the image type Img), and $o_{u}^{*}$ and $o_{i}^{*}$ are the projected representations of user $u$ and item $i$ in terms of * feature type, respectively. Using $o_{u}^{*}$ and $o_{i}^{*}$, we can then compute the binary cross-entropy loss $\mathcal{L}_{\text{RERI}}^{*}$ following Equation 1.
Among various item feature types, we extract the ID-based item embeddings from a pretrained collaborative filtering (CF) model (i.e., SASRec (Kang and McAuley, 2018)), thanks to its effectiveness in general recommendation tasks, especially for warm items (Kim et al., 2024b) [^7] In Section 4.4, we show that the utilization of ID-based item embeddings leads to more recommendations of warm items.. Specifically, we incorporate ID-based item embeddings (i.e., $c_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{CF}}$) in addition to the visual features (i.e., $v_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{Img}}$). Furthermore, as item descriptions are readily available in real-world scenarios (Zheng et al., 2024b), we can easily incorporate textual features $t$ extracted from full descriptions to further improve overall performance (i.e., $t_{i_{|S_{u}|}}=\mathbf{IF}^{\textsf{Text}}$). The analysis of extension to various features is provided in Section 4.4.
3.4 Training and Inference
Training. With the image, CF, and text feature types, we combine the training objective from RISA module and three training objectives from RERI module, training the adaptor $M$ and projectors $F^{*}_{u}$ and $F^{*}_{i}$ ($\textsf{Img},\textsf{CF},\textsf{Text}]$) while freezing the LLM. The combined loss [^8] We opt to fix all weights at 1 to avoid the computational burden of tuning over a large hyperparamter space. is denoted as:
$\begin{split}\mathcal{L}_{final}=\mathcal{L}_{\text{RISA}}+\mathcal{L}_{\text{RERI}}^{\textsf{Img}}+\mathcal{L}_{\text{RERI}}^{\textsf{CF}}+\mathcal{L}_{\text{RERI}}^{\textsf{Text}}\end{split}$ (2)
Inference. For retrieval-based inference, we compute the affinity scores $r_{u,i}^{\textsf{Img}}$, $r_{u,i}^{\textsf{CF}}$, and $r_{u,i}^{\textsf{Text}}$ using the scoring function $\mathcal{T}$, in terms of visual, CF, and textual features, respectively. Specifically, the top-$k$ relevant items for $i_{|\mathcal{S}_{u}|+1}$ is computed as $rec_{u}^{k}=\textsf{Top-k}(r_{u,i}^{\textsf{Img}}+r_{u,i}^{\textsf{CF}}+r_{u,i}^{\textsf{Text}}),\forall i\in\mathcal{I}$ , where $rec_{u}^{k}$ is the set of recommended $k$ items for user $u$, and Top-k is a function that extracts items with the highest top-$k$ scores. Note that while the affinity scores across different features can be aggregated through various approaches, such as weight summation or product, we opt for a simpler summation.
4 Experiments
  • RQ3: How does each module (i.e., RISA and RERI) contribute to the I-LLMRec?
  • RQ4: How can we handle when item images are missing?
4.1 Experimental Setting
Datasets. For evaluation, we use four categories from the Amazon dataset (Ni et al., 2019): Sports, Grocery, Art, and Phone. These datasets contain the item titles, attributes (e.g., brand and category), detailed textual descriptions, and images representing items. Following a prior study (Wei et al., 2024), we filter out users and items with fewer than 5 interactions to ensure data quality. We summarize the data statistics in the Appendix E.1.
Regarding the Implementation Details and Baselines, please refer to Appendix E.2 and E.3.
Table 1: Performance Comparison. $\mathbf{A}$: Attributed-based Representation, $\mathbf{CF}$: CF-based Representation (i.e., the CF item embedding is projected into the LLM), $\mathbf{D}$: Description-based Representation, $\mathbf{I}$: Image-based Representation.
Dataset Metric Collaborative Filtering LLM Image-aware LLM